# Exceptional Points in a $\mathcal{PT}$-symmetrical quantum system: a Scattering matrix approach

### 05 Dec 2023 v1

@@ Abstract
We analyze the behavior of a non-Hermitian opened one-dimensional quantum system with Parity-Time ($\mathcal{PT}$) symmetry. This system is built by a dimer, which has balanced gains and losses described by a parameter $\gamma$. By varying $\gamma$ the system resonances, which are naturally separated, coalesce at the exceptional point (EP). The transmission spectrum is obteined by means of the scattering matrix ($S$ matrix) formalism and we examine the wave functions corresponding to the resonances as a function of $\gamma$. Specifically, we look for the behavior and distribution of the phases of the S matrix before, at and after the exceptional point.
I Introduction
@@
Bender proposed in 1998 a new subset of non-Hermitian systems in which the energy spectrum has a real part Bender1998 . The Hamiltonians of this subset, which is called quasi-Hermitian, have the peculiarity of being invariant under the application of the symmetry operators of Parity, $\mathcal{P}$, and Time, $\mathcal{T}$. Also, regardless of whether this Hamiltonian is real or complex it has real eigenvalues Bender2002 ; Bender2007 . Systems with $\mathcal{PT}$-symmetry have a parameter associated with non-hermiticity that allows the transition between a real spectrum and a complex one. This transition was called Exceptional Point (EP) by Kato in 1966 Kato . EPs appear when the variation of this parameter allows the coalescence of two naturally repelling eigenvalues, as well as their eigenvectors, of the associated Hamiltonian operator  Alu2021 ; Heiss .
EPs arise when by varying a parameter associated with non-hermiticity, let us say $\gamma$, two or more eigenvalues coalesce. An example of a system with $\mathcal{PT}$-symmetry is a dimer made up of two coupled identical oscillators in which one of them has losses ($-\textrm{i}\gamma$) and the other one has gains ($+\textrm{i}\gamma$). The region prior to the critical value of the associated parameter ($0\leq\gamma<\gamma_{EP}$) is known as the exact $\mathcal{PT}$-phase. In this region, the eigenvalues only have a real part. When the associated parameter reaches the critical value $\gamma_{EP}$, the eigenvalues and eigenvectors coalesce in EP. Exceeding the EP we arrive at a region called broken $\mathcal{PT}$-phase in which the imaginary part of the eigenvalue is different from zero. In the exact $\mathcal{PT}$-phase the system behaves as an Hermitian one, while in the broken $\mathcal{PT}$-phase the system reveals its non-Hermitian nature. EPs have been found in various areas of physics such as electromagnetism electromagnetism2020 , optics optics2017 , acoustics acoustics2016 , photonics photonics2019 , elasticity vdomEP , quantum mechanics quantummechanics2018 , electronics electronics2017 or atomic physics atomicphysics2007 . EPs have been applied in improving the sensitivity of optical measurements optics2017 , in robust transfer of energy wireles electronics2017 , in unidirectional invisibility atomicphysics2007 , in alteration of spectral bands Zhen2015 ; Kartashov2018 , or in accelerometers KottosNat .
The system we study is made up of an open quantum dimer $\mathcal{PT}$-symmetric in one dimension. One of the scatterers of the dimer has losses while the other one has gains. Both losses and gains are balanced through a parameter associated with non-hermiticity. To study this system we use the formalism of the scattering matrix, or $S$ matrix. In essence we will study the transmission spectrum of the system, the phase of the $S$ matrix, its distribution, and the behavior of the wave function at resonances. These features will be analyzed as a function of $\gamma$ before, at, and after the EP.
The following section presents the system under study as well as the analityc expressions for the $S$ matrix. When varing the $\gamma$ parameter an EP appears in the transmission spectrum. Thus we show the evoution of the real and imaginary parts of the wavefunctions associated to the coalescing resonances. We study the behaviour of the phase, and its distribution, in the subsequent section. Conclusions are presented at the end of this paper.
II The $S$ matrix of a system with gain and losses
@@
The system we are studying is a one-dimensional open quantum dimer with balanced gains and losses to which plane waves are incident from both sides. The dimer is made up of two scatterers whose real part of the potential that describes them is identical, while the signs of the imaginary part are opposite to each other. Each scatterer of the dimer is composed of three potential barriers as illustrated in Figure 1. The potential that describes the whole system is composed by seven regions. It is real for all regions except for regions III and V where the potential is complex. In regions I and VII the potential $V(x)=0$; $V(x)=V_{1}$ for regions II (width $c$), IV (width $b$) and VI (width $c$); while the potential takes values $V_{2}-\textrm{i}\gamma$ for region III (width $a$) (losses), and $V_{2}+\textrm{i}\gamma$ for region V (width $a$) (gains). The width of the region IV is the sum of the barriers of each dimer scatterer since they are contiguous. For each of these regions the stationary solutions of the Schrödinger equation gives a superposition of plane waves traveling to the left and rigth given by $\psi_{1}(x)=a_{1}\textrm{e}^{\textrm{i}k_{1}x}+b_{1}\textrm{e}^{-\textrm{i}k_{1}x}$, and $\psi_{i}(x)=b_{i}\textrm{e}^{\textrm{i}k_{i}x}+a_{i}\textrm{e}^{-\textrm{i}k_{i}x}$ for $i=2,3,\cdots,7$. The wavenumbers are $k_{1}=k_{7}=\sqrt{2mE/\hbar^{2}}$, $k_{2}=k_{4}=k_{6}=\sqrt{2m(E-V_{1})/\hbar^{2}}$, $k_{3}=\sqrt{2m\left[E-(V_{2}-\textrm{i}\gamma)\right]/\hbar^{2}}$, and $k_{5}=\sqrt{2m\left[E-(V_{2}+\textrm{i}\gamma)\right]/\hbar^{2}}$, where $m$ is the mass of the particle, while $E$ is its energy ($V_{1}>E>V_{2}$).
@@
Figure 1: Diagram of a dimer composed of two scatterers, with three potential barriers each one. The potential of the outer barriers of each scatterer is real, while the potential described by the central barrier is complex. The sign of the imaginary part of the central barrier gives rise to losses (represented in the blue region) or gains (represented in the red region). As expected, this system is quasi-Hermitian since it is invariant when applying the operators $\mathcal{PT}$.
@@
The scattering matrix relates amplitudes of the outgoing waves in terms of the incoming ones PereyraBook ; MelloBook . In this case, the system has two opened channels FerryBook , thus the $S$ matrix is of $2\times 2$ and relates the amplitudes $b_{1}$ and $b_{7}$ in terms of the amplitudes $a_{1}$ and $a_{7}$, namely
@@
In former equations we have used two different notations explained at the Appendix. The Hermitian case is recovered when $\gamma=0$, thus $r=r^{\prime}$ and $t=t^{\prime}$ since the system has specular reflection and time reversal symmetry.
@@
Figure 2: Transmission spectra as a function of energy in which only the first doublet of the dimer is plotted. (a) For the Hermitian case ($\Gamma=0$) the maximum values of the resonances are found at $E_{1,1}=0.2086~{}\text{eV}$, where the first subscript refers to the associated doublet while the second one refers to the associated resonance, and $E_{1,2}=0.2615~{}\text{eV}$. Dimer parameters, in this and subsequent figures, are $a=1.15~{}$nm, $b=2c=0.02~{}\text{nm}$, $V_{1}=50~{}\text{eV}$, and $V_{2}=0~{}\text{eV}$. For $\Gamma=0$ the resonances have a unit amplitude and a maximum separation. As we increase the value of $\Gamma$ the separation between the resonances decreases and their amplitude increases before the EP as it is observed for $\Gamma$ equal to (b) 0.010, (c) 0.020. (d) 0.026 and (e) 0.0264. For $\Gamma=0.0264$ the resonances are already very close and overlap, losing the higher energy resonance. After the EP ($\Gamma\approx 0.0266$) the width of the resonance increases and its amplitude decreases as observed in panel (f). The same values of $\Gamma$ are used in Figures 3, 4, 7, 8, and 9.
@@
In Figure 2 the first doublet resonance spectra of the dimer is plotted as a function of energy. The resonances positions are obtained from the maxima of the total transmission ($T=|t|^{2}$). The Hermitian case is recovered for $\Gamma=0$ (see panel (a)), which is related with $\gamma$ through
$\Gamma=\sqrt{\left[\frac{\hbar^{2}}{m}\gamma^{2}+(E-V_{2})\right]^{2}-(E-V_{2})^{2}},$
@@
and the real and immaginary part of the wavefunctions of these resonances are plotted in panel (a) of Figures 3 and 4. All the next figures are obtained when incoming waves are only from the left, i.e. $a_{7}=0$. As can be anticipated, the real part of the wavefunction associated to the lowest energy resonance of the doublet is a symmetric one, while the highest energy resonance is antisymmetric. On the one hand, as we increase the $\Gamma$ value, the separation between the resonances decreases until they coalesce, as well as the width of the resulting resonance increases (see panels (b), (c), (d) and (e) of Figure 2). Also, the amplitude of the resonances increases but after passing the EP it starts decreasing its height (see Figure 2 (f)). This is notable since the amplitude of the resulting resonance is expected to grow indefinitely. On the other hand, as we increase $\Gamma$, the amplitude of the real and imaginary part of the wavefunction also increases. As can be seen in panels (b)-(e) of Figures 3 and 4, the amplitudes of the symmetric and antisymmetric modes grow although that of the symmetric mode grows more than the antisymmetric one. After passing the EP, in addition to the broken $\mathcal{PT}$-symmetry, both modes have coalesced into a mode that is neither symmetric nor antisymmetric. In fact, both modes coexist simultaneously and one is the complex conjugate of the other one. Figure 5 shows the behavior of the real and imaginary part of the energy as a function of $\Gamma$. The real part is associated with the position of the maxima of each resonance, while the imaginary part is associated with their width. The result given by the Tight Binding model predicts a decay in the separation between the eigenenergies of the system as a square root Kato ; Heiss . In Figure 6 it is shown the fit of the difference of the resonances as a function of $\Gamma$. In purple solid line the separation between the real parts of the resonances is plotted; the yellow solid line is a pure square root while the green solid line is the best fit that was obtained from data showing a power of 0.47.
@@
Figure 3: Evolution of the real part (Re$[\psi]$) of the states asociated to the resonances of the first doublet energy as the parameter $\Gamma$ is increased. The lowest energy resonance is plotted in blue while the one with highest energy is plotted in red, panels (a) to (e). The blue wavefunctions correspond to the symmetric state while the red ones are the antisymmetric states. Once the EP is reached only one resonance remains, therefore there is only one wave function as is shown in purple in panel (f). This mode is not symmetric nor antisymmetric. Before reaching the EP, the amplitude of the wave functions also increases until it decreases once the EP is passed.
@@
Figure 4: Evolution of the imaginary parts of the wavefunctions (Im$[\psi]$) associated to the resonances of Figure 2. The blue, red, curve is the imaginary part of the lowest, highest, energy resonance. As in the real part, also the amplitudes increase as a function of $\Gamma$. Once the EP is passed there is only one resonance and the amplitude of the imaginary part of its wavefunction decreases as well as the real part.
@@
Figure 5: EP of the dimmer. On the one hand, the real part corresponds to the positions of the resonances. For the Hermitian case, the separation of the resonances is given by the natural level repulsion. As $\Gamma$ increases, the separation between the resonances decreases until both coalesce at the EP. After the EP there is only one resonance since both have coalesced. On the other hand, the imaginary part corresponds to the widths of the resonances, measured by means of the $Q$ factor. As we get closer to the EP, the resonances overlap increasing their widths until they coalesce into a single one increasing significantly its width.
@@
Figure 6: The separation of the resonances as a function of $\Gamma$ is shown in purple. This is compared with the fit obtained, in green, which corresponds to the model $y=Ax^{B}$, where $A=0.061$ and $B=0.47$, the latter being the power obtained that is close to the square root predicted by the tight binding model, plotted in yellow.
III Phase distribution of the Scattering Matrix
Now let us study the real and imaginary part of the phase as a function of the energy, the phase in Argand plane, and the phase distribution as a function of $\Gamma$.
@@ III Phase distribution of the Scattering Matrix
The scattering matrix formalism allows to us to study not only the transmission and reflection amplitudes but also the phases of the system. The eigenphases of the $S$ matrix, given after diagonalizing Eq. 1Mares2009 , are
@@
In general the eigenphases are complex and can be written as
$\theta=\theta_{\text{Re}}+\textrm{i}\;\theta_{\text{Im}},$ (11)
@@
In Figure 7 we plot the real and the imaginay part of the phase as a function of the energy. Panel (a) corresponds to the Hermitian case in wich the real part has two separated resonances that correspond to $\theta$ and $\theta^{\prime}$. In this case the imaginary part is null. As $\Gamma$ increases the distance between resonances becomes smaller but the imaginary part still been zero as can be seen in panels (b) and (c). In panel (d) the resonances overlap and the imaginary part becomes not null for positive and negative values. This protuberance appears near the lowest energy resonance, which is the one that prevails. As we approach the EP the separation of the resonances reduces while the protuberance grows (panel (e)). Once the EP value has been exceeded the real part of the phase is restricted to values close to $-\pi$ and $\pi$ and the imaginary part remains not null in the neighborhood of the remain resonance as is plotted in panel (f).
Figure 7: Behaviour of the real (yellow solid line) and imaginary (brown solid line) parts of the phases $\theta$ and $\theta^{\prime}$ as a function of the energy. (a) In the Hermitian case the separation of the resonances is maximum and the imaginary part of the phase is zero. As $\Gamma$ grows the separation of the resonances diminishes (panels (b) and (c)). (d) When we approach to the EP the imaginary part stops being null and a protuberance appears near the lowest energy resonance. (e) The closer we get to the EP the more the protuberance grows. (f) Once we past the EP, the real part of the phases lives only in the neighborhood of $-\pi$ and $\pi$, and the protuberance becomes smaller.
@@
Figure 8: Phase distribution of both resonances in Argand plane. Blue points corresponds to the lowest energy resonance while the red ones corresponds to the highest energy resonance. In (a), (b), and (c) all the points live in the unit circle because $\theta_{\text{Im}}=0$ (blue points are bellow the red ones). In panels (d) and (e) the red resonance still lives in the unit circle but a lobe appears in the lowest energy resonance. This lobe increases its size up to infinity where it changes its concavity to reduce its size as $\Gamma$ keeps growing. In panel (f) the coalesced resonance has diminished its size and distribution.
@@
Figure 9: Histogram of the phase distribution of the resonances as a function of $\Gamma$. Blue (red) histogram corresponds to the lowest (highest) energy resonance, except for the coaleced resonance plotted in purple. As can be inferred for Argand plane, the phases are distributed all over the angles but not in a uniform way as is shown in (a), (b) and (c). When the lobe appears the resonances stop visiting all the angles so empty spaces appear in the distribution (panels (d) and (e)). In (f) all of the poins are distributed only around $-\pi$ and $\pi$.
IV Conclusions
In this paper we have studied a quantum system in one dimension in which electrons, described by plane waves, had sent. In particular, the system that was studied consists of a dimer made up of two scatterers that presented losses and gains, parameterized by $\Gamma$. Electrons were incident on the dimer and this was described by means of the scattering matrix formalism. The $S$ matrix provided us with the transmission coefficient from which the resonances of the system were obtained. The transmission spectrum was studied as a function of $\Gamma$ and the coalescence of the resonances associated with the first doublet was observed, thus obtaining an Exceptional Point. It was observed that the separation between the resonances decays as a power as $\Gamma$ increases, and when making the fit we obtained that the power is basically a square root. Analytic expressions were developed for the elements of the scattering matrix in the non-Hermitian case.
Once the EP was obtained, we analyze the properties of the system such as the wave functions, the resonances, and the phase, before and after the EP. In the diagram of the phase against the energy we observe that before the EP the phases are separated and as we increase $\Gamma$ they join; close to the EP the real part of the phase breaks while the imaginary part presents a lobe that increases its size; after the EP the phase only exists in values close to $-\pi$ and $\pi$. For the Argand plane case before the EP, we observed that initially each resonance formed a unit circle; near the EP the circles began to have a redistribution of points, while a lobe arose. This lobe increased in size significantly near the EP and once reached, the size of the lobe decreased. All the numerical evidence that we analyze suggests a saturation process that comes from inducing gains in the system. In the phase histogram the distribution of points changed as we reached the EP. Initially the points were distributed in the entire range between $-\pi$ and $\pi$, and this changed drastically when we approached the EP, firstly, appearing gaps in the distribution and then increasing the size of the empty regions as we reach the EP. After the EP the points were redistributed only around $-\pi$ and $\pi$.

### 10 Apr 2024 v2

# The rise and fall of the amplitude, and phase, around Exceptional Points: a Scattering matrix approach

@@ Abstract
We analyze the behavior of a non-Hermitian opened one-dimensional quantum system with $\mathcal{PT}$ symmetry. This system is built by a dimer, with balanced gains and losses described by a parameter $\gamma$. By varying $\gamma$ the system resonances, which are naturally separated, coalesce at the exceptional point (EP). The transmission spectrum is obtained by means of the scattering matrix ($S$ matrix) formalism and we examine the wave functions corresponding to the resonances as a function of $\gamma$. Specifically, we look for the behavior and distribution of the phases of the $S$ matrix before, at and after the EP.
@@
A complete generalization of non-Hermitian quantum mechanics has not been reached yet since Bender proposed, more than two decades ago, a new subset of systems in which the energy spectrum has an imaginary part Bender1998 . The Hamiltonians of this subset, which are called quasi-Hermitian, have the peculiarity of being invariant under the application of the symmetry operators of parity (${\cal P}$) and time (${\cal T}$). Also, regardless of whether this Hamiltonian is real or complex it has real eigenvalues Bender2002 ; Bender2007 . Systems with ${\cal PT}$-symmetry have a parameter associated with non-hermiticity that allows the transition between a real spectrum and a complex one. This special value was called Exceptional Point (EP) by Kato in 1966 Kato . EPs appear when the variation of that parameter allows the coalescence of two (or more) naturally repelling eigenvalues, as well as their eigenvectors, of the associated Hamiltonian operator  Alu2021 ; Heiss . Quasi-Hermitian systems and EPs have been extensively studied in various areas of physics such as electromagnetism electromagnetism2020 , optics optics2017 ; Chong2011 ; Ramezani2011 , microwaves Richter , acoustics acoustics2016 , photonics photonics2019 , elasticity vdomEP , quantum mechanics Schomerus ; quantummechanics2018 ; Ortega ; Dorey , electronics electronics2017 or atomic physics atomicphysics2007 . EPs have been applied in improving the sensitivity of optical measurements optics2017 , in robust transfer of energy wireless electronics2017 , in unidirectional invisibility Chong2010 ; atomicphysics2007 , slowing light Moiseyev2018 , in alteration of spectral bands Zhen2015 ; Kartashov2018 , or in accelerometers KottosNat .
In order to try to bring some additional information to the global discussion, in this letter we wonder how the distribution of the phases of a scattering system changes as a function of the non-hermicity parameter $\gamma$. To do this, we study the properties of an open one-dimensional non-Hermitian quantum system which presents at least one EP. The system is made up of a ${\cal PT}$-symmetric dimer, where one of the scatterers has losses while the other one has gains. Both losses and gains are balanced through $\gamma$. We study the transmission spectra by developing analytical expressions for the scattering matrix ($S$ matrix) of the dimer. We show that the transmission increases its amplitude as $\gamma$ grows in the exact ${\cal PT}$-phase but it starts decreasing its amplitude as we move further into the broken ${\cal PT}$-phase. In fact, the $S$ matrix tells to us about the unitarity of the system. On the one hand, it is quite common to think that a system has absorption if $S$ is subunitary ($R+T<1$). On the other hand, we must think that a system is emiting if $S$ is over unitary Schomerus ($R+T>1$). The system we are studying has both behaviours. In the process of showing this behaviour, we analyze the wave functions of the resonances involved in the EP. The phases of the $S$ matrix in the Argand plane support this behaviour by increasing its radii before and after the EP. Even more surprising is the fact that the distribution of the phases cluster around a certain angle when they were distributed all over the Argand plane.
@@
The system we are studying is a one-dimensional open quantum dimer with balanced gains and losses, parameterized by $\gamma$, to which plane waves are incident from both sides. The dimer is made up of two scatterers whose real part of the potential that describes them is identical, while the signs of the imaginary part are opposite to each other. Each scatterer of the dimer is composed of three potential barriers as illustrated in Figure 1. The potential that describes the whole system is composed by seven regions. It is real for all regions except for regions III and V where the potential is complex. In regions I and VII the potential $V(x)=0$; $V(x)=V_{\textrm{b}}$ for regions II, IV and VI; while the loss (gain) potential is $V^{\prime}-\textrm{i}\gamma$ ($V^{\prime}+\textrm{i}\gamma$) for region III (V). The width of the region IV is the sum of the barriers of each dimer scatterer since they are contiguous. For each of these regions the stationary solutions of the Schrödinger equation gives a superposition of plane waves traveling to the left and right given by $\psi_{1}(x)=a_{1}\textrm{e}^{\textrm{i}k_{1}x}+b_{1}\textrm{e}^{-\textrm{i}k_{1}x}$, and $\psi_{i}(x)=b_{i}\textrm{e}^{\textrm{i}k_{i}x}+a_{i}\textrm{e}^{-\textrm{i}k_{i}x}$ for $i=2,3,\cdots,7$. The wavenumbers are $k_{1}=k_{7}=\sqrt{2mE/\hbar^{2}}$, $k_{2}=k_{4}=k_{6}=\sqrt{2m(E-V_{\textrm{b}})/\hbar^{2}}$, $k_{3}=\sqrt{2m\left[E-(V^{\prime}-\textrm{i}\gamma)\right]/\hbar^{2}}$, and $k_{5}=\sqrt{2m\left[E-(V^{\prime}+\textrm{i}\gamma)\right]/\hbar^{2}}$, where $m$ is the mass of the particle, while $E$ is its energy. We are only interested in resonances, consequently the energy must be bounded between the heights of the barriers as $V_{\textrm{b}}>E>V^{\prime}$.
We are only interested in resonances, therefore the energy must be $V_{\textrm{b}}>E>V^{\prime}$. We restrict ourselves to the energy region $V_{\textrm{b}}>E>V^{\prime}$ because we are focus on the resonance phenomena It could be interesting to study this phenomenon in the region below We restrict ourselves to the energy region since we are not interested in the evanescent case
Resonance phenomena
@@
Figure 1: (Color online) Diagram of a dimer composed of two scatterers, with three potential barriers each one. The potential of the outer barriers of each scatterer is real, while the potential described by the central barrier is complex. The sign of the imaginary part of the central barrier gives rise to losses (blue region) or gains (red region). As expected, this system is quasi-Hermitian since it is invariant when applying the operators ${\cal PT}$.
@@
The scattering matrix relates the amplitudes of the outgoing waves ($b_{1}$ and $b_{7}$) in terms of the incoming ones ($a_{1}$ and $a_{7}$ereyraBook ; MelloBook . The $S$ matrix is of $2\times 2$ because the system has two opened channels FerryBook , namely
@@
In former equations we have used two different notations explained at the Appendix. The Hermitian case is recovered when $\gamma=0$, therefore $r=r^{\prime}$ and $t=t^{\prime}$ since the system has specular reflection and time reversal symmetry. We must emphasize that these equations are analytical and general expressions regardless of the nature of the waves as long as the definition of the scattering matrix for one-dimensional cases must be fulfilled Robledo2020 .
@@
Figure 2: (Color online) Transmission spectra as a function of energy in which only the first doublet of the dimer is plotted. (a) For the Hermitian case ($\Gamma=0$) the maximum values of the resonances are found at $E_{1,1}=0.2086~{}\text{eV}$ (the first subscript refers to the associated doublet while the second one refers to the associated resonance), and $E_{1,2}=0.2615~{}\text{eV}$. Dimer parameters, in this and subsequent figures, are $a=1.15~{}$nm, $b=2c=0.02~{}\text{nm}$, $V_{\textrm{b}}=50~{}\text{eV}$, and $V^{\prime}=0~{}\text{eV}$. For $\Gamma=0$ the resonances have a unit amplitude and a maximum separation. As we increase the value of $\Gamma$ the separation between the resonances decreases and their amplitude increases before the EP as it is observed for $\Gamma$ equal to (b) 0.010, (c) 0.020. (d) 0.026 and (e) 0.0264. For $\Gamma=0.0264$ the resonances are already very close and overlap, losing the higher energy resonance. After the EP ($\Gamma\approx 0.0266$) the width of the resonance increases and its amplitude decreases as observed in panel (f). The same values of $\Gamma$ are used in Figures 3, 4, 7, 8, and 9.
@@
In Figure 2 the first doublet of the dimer resonance spectrum is plotted as a function of energy. The resonances positions are obtained from the maxima of the total transmission ($T=|t|^{2}$). The Hermitian case is recovered for $\Gamma=0$ (see Figure 2(a)), which is related with $\gamma$ through
$\Gamma=\sqrt{\left[\frac{\hbar^{2}}{m}\gamma^{2}+(E-V^{\prime})\right]^{2}-(E-V^{\prime})^{2}},$
@@
and the real and imaginary parts of the wavefunctions of these resonances are plotted in panel (a) of Figures 3 and 4, respectively. All the next figures are obtained when incoming waves are only from the left, i.e. $a_{7}=0$. As can be anticipated, the real part of the wavefunction associated to the lowest energy resonance of the doublet is a symmetric one, while the highest energy resonance wavefunction is antisymmetric. On the one hand, as we increase the $\Gamma$ value, the separation between the resonances decreases until they coalesce, as well as the width of the resulting resonance increases (see panels (b), (c), (d) and (e) of Figure 2). Also, the amplitude of the resonances increases but after passing the EP it starts decreasing its height (see Figure 2 (f)). This is outstanding because the amplitude of the resulting resonance is expected to grow indefinitely. On the other hand, as we increase $\Gamma$, the amplitude of the real and imaginary part of the wavefunction also increases. As can be seen in panels (b)-(e) of Figures 3 and 4, the amplitudes of the symmetric and antisymmetric modes grow although that of the symmetric mode grows more than the antisymmetric one. After passing the EP, in addition to the broken $\mathcal{PT}$-symmetry, both modes have coalesced into a mode that is neither symmetric nor antisymmetric. In fact, both modes coexist simultaneously and one is the complex conjugate of the other one. Figure 5 shows the behavior of the real and imaginary part of the energy as a function of $\Gamma$. The real part is associated with the position of the maxima of each resonance, while the imaginary part is associated with their width. The result given by the Tight Binding model predicts a decay in the separation between the eigenenergies of the system as a square root Kato ; Heiss . In Figure 6 it is shown the fit of the difference of the resonances as a function of $\Gamma$. In purple solid line the separation between the real parts of the resonances is plotted; the yellow solid line is a pure square root while the green solid line is the best fit that was obtained from data showing a power of 0.47.
@@
Figure 3: (Color online) Evolution of the real part (Re$[\psi]$) of the states associated to the resonances of the first doublet energy as the parameter $\Gamma$ is increased. The lowest energy resonance is plotted in blue while the one with highest energy is plotted in red, panels (a) to (e). The blue wavefunctions correspond to the symmetric state while the red ones are the antisymmetric states. Once the EP is reached only one resonance remains, therefore there is only one wave function as is shown in purple in panel (f). This mode is not symmetric nor antisymmetric. Before reaching the EP, the amplitude of the wave functions also increases until it decreases once the EP is passed.
@@
Figure 4: (Color online) Evolution of the imaginary parts of the wavefunctions (Im$[\psi]$) associated to the resonances of Figure 2. The blue, red, curve is the imaginary part of the lowest, highest, energy resonance. As in the real part, also the amplitudes increase as a function of $\Gamma$. Once the EP is passed there is only one resonance and the amplitude of the imaginary part of its wavefunction decreases as well as the real part.
@@
Figure 5: (Color online) EP of the dimmer. On the one hand, the real part corresponds to the positions of the resonances. For the Hermitian case, the separation of the resonances is given by the natural level repulsion. As $\Gamma$ increases, the separation between the resonances decreases until both coalesce at the EP. After the EP there is only one resonance since both have coalesced. On the other hand, the imaginary part corresponds to the widths of the resonances, measured by means of the $Q$ factor. As we get closer to the EP, the resonances overlap increasing their widths until they coalesce into a single one increasing significantly its width.
@@
Figure 6: (Color online) The separation of the resonances as a function of $\Gamma$ is shown in purple. This is compared with the obtained fit, in green, which corresponds to the model $y=Ax^{B}$, where $A=0.061$ and $B=0.47$, the latter being the power obtained that is close to the square root predicted by the tight binding model, plotted in yellow.
@@
The scattering matrix formalism allows to us to study not only the transmission and reflection amplitudes but also the phases of the system. The eigenphases of the $S$ matrix, given after diagonalizing Eq. 1, are
@@
For $\Gamma=0$ also the Hermitian eigenphases are recovered Mares2009 . In general, the eigenphases are complex and can be written as
$\theta=\theta_{\text{Re}}+\textrm{i}\;\theta_{\text{Im}},$
@@
In Figure 7 we plot the real and the imaginary part of the phase as a function of the energy. Panel (a) corresponds to the Hermitian case in wich the real part has two separated resonances that correspond to $\theta$ and $\theta^{\prime}$. In this case the imaginary part is null. As $\Gamma$ increases the distance between resonances becomes smaller but the imaginary part still been zero as can be seen in panels (b) and (c). In panel (d) the resonances overlap and the imaginary part becomes not null for positive and negative values. This protuberance appears near the lowest energy resonance, which is the one that prevails. As we approach to the EP the separation of the resonances reduces while the protuberance grows (panel (e)). Once the EP value has been exceeded the real part of the phase is restricted to values close to $-\pi$ and $\pi$ and the imaginary part remains not null in the neighborhood of the remain resonance as is plotted in panel (f).
Figure 7: (Color online) Behaviour of the real (yellow solid line) and imaginary (brown solid line) parts of the phases $\theta$ and $\theta^{\prime}$ as a function of the energy. (a) In the Hermitian case the separation of the resonances is maximum and the imaginary part of the phase is zero. As $\Gamma$ grows the separation of the resonances diminishes (panels (b) and (c)). (d) When we approach to the EP the imaginary part stops being null and a protuberance appears near the lowest energy resonance. (e) The closer we get to the EP the more the protuberance grows. (f) Once we past the EP, the real part of the phases lives only in the neighborhood of $-\pi$ and $\pi$, and the protuberance becomes smaller.
@@
Figure 8: (Color online) Phase distribution of both resonances in Argand plane. Blue points correspond to the lowest energy resonance while the red ones correspond to the highest energy resonance. In (a), (b), and (c) all the points live in the unit circle because $\theta_{\text{Im}}=0$ (blue points are below the red ones). In panels (d) and (e) the red resonance still lives in the unit circle but a lobe appears in the lowest energy resonance. This lobe increases its size up to infinity where it changes its concavity to reduce its size as $\Gamma$ keeps growing. In panel (f) the coalesced resonance has diminished its size and distribution.
@@
Figure 9: (Color online) Histogram of the phase distribution of the resonances as a function of $\Gamma$. Blue (red) histogram corresponds to the lowest (highest) energy resonance, except for the coalesced resonance plotted in purple. As can be inferred from the Argand plane, the phases are distributed all over the angles but not in a uniform way as is shown in (a), (b) and (c). When the lobe appears the resonances stop visiting all of the angles hence empty spaces appear in the distribution (panels (d) and (e)). In (f) all of the points are distributed only around $-\pi$ and $\pi$.
I Concluding remarks
We have shown that the amplitude of the resonances of a ${\cal PT}$-symmetric system which presents an EP grows as $\Gamma$ get closer to the transition value. At the EP the resonances coalesce in energy as well as the eigenstates. The amplitude of the remain resonance starts decreasing as $\Gamma$ keeps growing. The eigenphases show the same behaviour by increasing their radii up to infinity and becoming smaller once we have past the EP. The distribution of the phase changes dramatically starting from being distributed in all over the angles to be localized in the vicinity of a well defined region. Since we have developed general expressions for the scattering matrix these behaviours can be tested experimentally in microwaves, or elastic, regime by building a proper system with slabs, or notches, outside the gain and loss regions in the one-dimensional approximation. Our desire that the results presented in this letter will contribute to the generalization of a non-Hermitian quantum theory.